A firm is producing a product using only labor and physical capital. The price of labor (wage rate) is $ 20/hour, and the cost of capital is $ 60/hour. If the marginal product of labor is 300 units an hour, what must be the marginal product of capital if the firm is a profit maximizer?      
Answer #1

As it has been given that if A firm is producing a product using only labor and physical capital. The price of labor (wage rate) is $ 20/hour, and the cost of capital is $ 60/hour. If the marginal product of labor is 300 units an hour.

If the firm is a profit maximizer, then following condition must satisfy

MPL/MPK=PL/PK

300/MPK=20/60

300/MPK=1/3

MPK=300*3

=900 units an hour.

Hence the Marginal product of capital must be 900 units an hour.